# An Empirical Risk Minimization Approach for Offline Inverse RL and Dynamic Discrete Choice Model

### v4

@@ 1 Introduction
Our contributions. In this paper, we propose an Empirical Risk Minimization (ERM)–based gradient-based method for IRL/DDC as an inverse Q-learning method. This method provably finds the true parameter for $Q^{\ast}$ estimation (up to statistical error, which diminishes at an $O(1/\sqrt{N})$ rate with $N$ samples) with $O(1/\sqrt{T})$ rate of convergence, where $T$ is the number of gradient iterations. In addition, the true reward function can be computed from the estimated $Q^{\ast}$ with no extra statistical or computational cost, given the estimated $Q^{\ast}$ function. In developing this method, we make the following technical contributions:
  • We propose an empirical risk minimization (ERM) problem formulation, which we refer to as ERM-IRL in the IRL literature and ERM-DDC in the DDC literature, reflecting the shared problem. This formulation allows us to circumvent the need for explicit transition function estimation. [^2] Transition function estimation can also be avoided in the counterfactual reasoning stage. Once the estimated reward function $\hat{r}$ is in hand, one can perform counterfactual policy evaluation and optimization by running an offline RL routine over a reward info-augmented data of $(s,a,\hat{r})$ tuples. Notably, this formulation also allows us to conclude that imitation learning (IL) is a strictly easier problem than IRL/DDC estimation problem.
@@ 2.1 Dynamic discrete choice model estimation literature
Adusumilli and Eckardt (2019) proposed how to adapt two popular temporal difference (TD)-based methods (an approximate dynamic programming-based method and a semi-gradient descent method based on Tsitsiklis and Van Roy (1996a)) for DDC. As discussed earlier, approximate dynamic programming-based methods are known to suffer from a lack of provable convergence beyond linear reward models (Jiang and Xie 2024; Tsitsiklis and Van Roy 1996b; Van Hasselt et al. 2018; Wang et al. 2021) [^4] In fact, “fitted fixed point methods can diverge even when all of the following hold: infinite data, perfect optimization with infinite computation, 1-dimensional function class (e.g., linear) with realizability (Assumption 4)(the instability is) not merely a theoretical construction: deep RL algorithms are known for their instability and training divergence…(Jiang and Xie 2024); the semi-gradient method is a popular, efficient approximation method that has theoretical assurance of convergence to projected squared Bellman error minimizers for linear reward/value functions (Sutton and Barto 2018). Feng et al. (2020) showed global concavity of value function under certain transition functions and monotonicity of value functions in terms of one-dimensional state, both of which are easily satisfied for applications in social science problems. However, those conditions are limitedly satisfied for the problems with larger dimensional state space.
@@ 2.2 Offline inverse reinforcement learning literature
The most widely used inverse reinforcement learning model, Maximum-Entropy inverse reinforcement learning (MaxEnt-IRL), assumes that the random choice happens due to agents choosing the optimal policy after penalization of the policy by its Shannon entropy (Ermon et al. 2015). In addition to the equivalence of MaxEnt-IRL to DDC (See Ermon et al. (2015) and Web Appendix §$\S$3.3), the identifiability condition for DDC (Magnac and Thesmar 2002) was rediscovered by Cao et al. (2021) for MaxEnt-IRL. Zeng et al. (2023) proposes a two-step maximum likelihood-based method that can be considered as a conservative version of CCP method of Hotz and Miller (1993) [^5] When there is no uncertainty in the transition function, approximated trajectory gradient of Offline IRL method degenerates to forward simulation-based gradient in CCP estimator method of Hotz and Miller (1993).. Despite that their method is proven to be convergent, it requires the explicit estimation of the transition function, and its global convergence was only proven for linear reward functions.
Finn et al. (2016) showed that a myopic [^6] See Cao et al. (2021) for more discussion on this. version of MaxEnt-IRL can be solved by the Generative Adversarial Network (GAN) training framework (Goodfellow et al. 2020). Fu et al. (2017) extended this framework to a non-myopic version, which is proven to recover the reward function and the value function under deterministic transitions. The GAN approach has been extended to $Q$-estimation methods that use fixed point iteration (Geng et al. 2020, 2023). Ni et al. (2021) has shown that the idea of training an adversarial network can also be used to calculate the gradient direction for minimizing the myopic version of negative log likelihood [^7] Minimizing negative log-likelihood is equivalent to minimizing KL divergence. See the Proof of Lemma 20.. Zhan et al. (2024) proposed a GAN approach that is provably equivalent to negative log likelihood minimization objective (i.e., behavioral cloning (Torabi et al. 2018)). Indeed, GAN approaches are known to work well for Imitation Learning (IL) tasks (Zare et al. 2024). However, as discussed earlier, solving for IL does not allow us to conduct a counterfactual simulation.
@@ 3.1 Setup: Maximum Entropy-Inverse Reinforcement Learning (MaxEnt-IRL)
Throughout the paper, we use $\lambda=1$, the setting which is equivalent to dynamic discrete choice (DDC) model with mean zero T1EV distribution (see Web Appendix §$\S$3.3 for details); all the results of this paper easily generalize to other values of $\lambda$. Given $\pi^{*}$, we define the value function $V^{*}$ as:
@@ 3.4 Identification
In Lemma 1, we formally establish that Assumptions 1 and 3 uniquely identify $Q^{\ast}$ and $r$. See Web Appendix §$\S$§C.2 for the proof.
@@
gives a unique $Q=Q^{*}$. Furthermore, $r$ is obtained by solving:
@@ 4.1 Identification via expected risk minimization
[…]
@@ 4.2 Estimation via minimax-formulated empirical risk minimization
Temporal Difference (TD) methods typically use stochastic approximation to obtain an estimate of this proxy objective (Tesauro et al. 1995; Adusumilli and Eckardt 2019). However, the issue with TD methods is that minimizing the proxy objective will not minimize the Bellman error in general (see Web Appendix §$\S$§C.1 for details), because of the extra variance term, as shown below.
@@
By further re-parametrizing $h$ using $\zeta=h-r(s,a)+Q(s,a)$, after some algebra, we arrive at Lemma 4. See Web Appendix §$\S$§C.1 for the detailed derivation.
@@
$\mathcal{R}_{emp}(Q;\mathcal{D}):=\frac{1}{N}\sum_{\left(s,a,s^{\prime}\right)\in\mathcal{D}}\bigl{[}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}-\log\left(\hat{p}_{Q}(a\mid s)\right)}\bigr{]}+\frac{1}{N}\sum_{\left(s,a,s^{\prime}\right)\in\mathcal{D}}\bigl{[}\mathbbm{1}_{a=a_{s}}\bigl{\{}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bigl{(}\hat{\mathcal{T}}Q\left(s,a,s^{\prime}\right)-Q(s,a)\bigr{)}^{2}}\bigr{\}}\bigr{]}$
$\quad\quad\quad\quad\quad\quad-\frac{1}{N}\sum_{\left(s,a,s^{\prime}\right)\in\mathcal{D}}\bigl{[}\mathbbm{1}_{a=a_{s}}\bigl{\{}\beta^{2}{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\left(V_{Q}\left(s^{\prime}\right)-\bar{\zeta}(s,a)\right)^{2}}\bigr{\}}\bigr{]}$ (15)
@@ 5 GLADIUS: Algorithm for ERM-IRL (ERM-DDC)
8:     where $V_{\bm{\theta}}(s^{\prime}):=\log\sum_{\tilde{a}\in\mathcal{A}}\exp(Q_{\bm{\theta}}(s^{\prime},\tilde{a}))$
@@
18:$\widehat{\zeta}\leftarrow\zeta_{\bm{\theta_{1}}}$
19:$\widehat{Q}\leftarrow Q_{\bm{\theta_{2}}}$
20:$\widehat{r}(s,a)\leftarrow\widehat{Q}(s,a)-\beta\cdot\widehat{\zeta}(s,a)$
@@
13:$\widehat{Q}\leftarrow Q_{\bm{\theta}}$
14:$\widehat{r}(s,a)\leftarrow\widehat{Q}(s,a)-\beta\log\sum_{\tilde{a}\in\mathcal{A}}\exp(\widehat{Q}(s^{\prime},\tilde{a}))$
When the transition function is deterministic (e.g., in Rafailov et al. (2024); Guo et al. (2025); Zhong et al. (2024)) meaning that for any state-action pair $(s,a)$, the next state $s^{\prime}$ is uniquely determined, the ascent step involving $\zeta$ is no longer required. This is because the term $\left(V_{Q}\left(s^{\prime}\right)-\zeta(s,a)\right)^{2}$ (highlighted in orange in Equation (12) and (15)) becomes redundant in the empirical ERM-IRL objective, because $\max_{\zeta\in\mathbb{R}}-\mathbb{E}_{s^{\prime}\sim P(s,a)}\left[\left(V_{Q}\left(s^{\prime}\right)-\zeta\right)^{2}\mid s,a\right]$ is always 00. Consequently, the optimization simplifies to:
@@ 6 Theory and analysis of GLADIUS
As discussed in the previous section, Equation (15) represents a mini-max optimization problem. Such problems are known to be globally solvable by a simple gradient ascent-descent algorithm when it is a concave-convex mini-max problem. However, the challenge is that Equation (12) is not a concave-convex mini-max problem. Given $Q$, it determines $\zeta$ that serves as the Bayes-optimal estimator for $\mathbb{E}_{s^{\prime}\sim P(s,a)}\left[V_{Q}\left(s^{\prime}\right)\mid s,a\right]$. This implies that $-\mathbb{E}_{s^{\prime}\sim P(s,a)}\left[\left(V_{Q}\left(s^{\prime}\right)-\zeta\right)^{2}\mid s,a\right]$ is strongly concave in $\zeta$. On the other hand, given such an optimal $\zeta$, $\mathcal{L}_{BE}(Q)(s,a)$ term is not convex in $Q$ (Bas-Serrano et al. 2021). The key result in this section is proving that both $\mathcal{L}_{BE}(Q)(s,a)$ and $\mathcal{L}_{\mathrm{NLL}}(Q)(s,a)=\left[-\log\left(\hat{p}_{Q}(a\mid s)\right)\right]$, under certain parametrization of $Q$, satisfies the Polyak-Łojasiewicz (PL) condition, which allows Algorithm 1 to converge to global optima.
@@ 6.2 Global convergence of GLADIUS
Denote $\mathcal{D}_{N}$ be a finite-size dataset with $N$ number of transition pairs $((s,a,s^{\prime}))$. We define $\bm{\theta}^{\ast}_{N}$ as the solution to the empirical risk minimization problem (minimizing Equation (18)), i.e.,
$\bm{\theta}^{\ast}_{\mathcal{D}_{N}}\in\arg\min_{\bm{\theta}\in\Theta}$ $\max_{\zeta\in\mathbb{R}^{S\times A}}\frac{1}{N}\sum_{(s,a,s^{\prime})\in\mathcal{D}_{N}}$
$\bigl{[}{-\log\left(\hat{p}_{Q_{\bm{\theta}}}(a\mid s)\right)}+\mathbbm{1}_{a=a_{s}}\bigl{\{}\bigl{(}\hat{\mathcal{T}}Q_{\bm{\theta}}\left(s,a,s^{\prime}\right)-Q_{\bm{\theta}}(s,a)\bigr{)}^{2}-\beta^{2}\bigl{(}{\left(V_{Q_{\bm{\theta}}}\left(s^{\prime}\right)-\zeta(s,a)\right)^{2}}\bigr{)}\bigr{\}}\bigr{]}$
Note that $\bm{\theta}^{\ast}_{\mathcal{D}_{N}}$ approaches to $\bm{\theta}^{\ast}$ as $N\rightarrow\infty$, the solution to the ERM-IRL objective (Equation (12)), by the law of large numbers. Formally, we define the statistical error, $\epsilon_{\text{stat}}(N)$, as
$\epsilon_{\text{stat}}(N):=\mathbb{E}_{\mathcal{D}_{N}}\left[\left(\bm{\theta}_{\mathcal{D}_{N}}^{*}-\bm{\theta}^{*}\right)^{2}\right]$
Also, define $\hat{\bm{\theta}}_{T}$ be the $T$ iteration outcome of Algorithm 1. Then we can define the optimization error of GLADIUS (Algorithm 1) at its $T$th iteration given data $\mathcal{D}$, $\epsilon_{\text{opt}}(T)$, as
$\epsilon_{\text{opt}}(T):=\left(\bm{\theta}^{\ast}_{\mathcal{D}_{N}}-\hat{\bm{\theta}}_{T}\right)^{2}$
@@
$\mathbb{E}_{\mathcal{D}_{N}}\left[\left(\theta^{*}-\hat{\theta}_{T}\right)^{2}\right]\leq 2\left[\epsilon_{\text{stat }}(N)+\mathbb{E}_{\mathcal{D}_{N}}\left[\epsilon_{\text{opt }}(T)\right]\right]$ $\leq\frac{2\nu}{\gamma+T}+\frac{2\sigma}{N}$
where the values of constants $\gamma$, $\nu$, and $\sigma$ are specified in the Web Appendix §$\S$§C.8 where we prove Theorem 11.
@@ 7 Offline IRL/DDC experiments
We now present results from simulation experiments, comparing the performance of our approach with that of several benchmark algorithms. In this section, we use the high-dimensional version of the canonical bus engine replacement problem (Rust (1994)) as the setting for our experiments. Later, in §$\S$8, we use the OpenAI gym benchmark environment experiments with a discrete action space (Lunar Lander, Acrobot, and Cartpole) Brockman (2016) as in Garg et al. (2021) for the related but easier problem of imitation learning.
@@ 7.1 Experimental Setup
State transitions at each period.   If the manager chooses maintenance, the mileage advances by 1, 2, 3, or 4 with a $1/4$ probability each. If the manager chooses to replace the engine, then the mileage is reset to 1. That is, ($\mathbb{P}(\{x_{j(h+1)}=x_{jh}+k\}\mid d_{jh}=0)=1/4$, $k\in\{1,2,3,4\}$ and $\mathbb{P}\{x_{j(h+1)}=1\mid d_{jh}=1\}=1$. When the bus reaches the maximum mileage of 20, we assume that mileage remains at 20 even if the manager continues to choose maintenance.

### v6

@@ 1 Introduction
Our contributions. In this paper, we propose an Empirical Risk Minimization (ERM)–based gradient-based method for IRL/DDC as an inverse Q-learning method. This method provably finds the true parameter $\bm{\theta}$ for $Q^{*}$ estimation (up to statistical error, which diminishes at an $O(N^{-1/2})$ rate with $N$ samples) with $O(T^{-1/4})$ rate of convergence, where $T$ is the number of gradient iterations. In addition, the true reward function can be computed from the estimated $Q^{\ast}$ with no extra statistical or computational cost, given the estimated $Q^{\ast}$ function. In developing this method, we make the following technical contributions:
[…]
@@ 2.1 Dynamic discrete choice model estimation literature
Adusumilli and Eckardt (2019) proposed how to adapt two popular temporal difference (TD)-based methods (an approximate dynamic programming-based method and a semi-gradient descent method based on Tsitsiklis and Van Roy (1996b)) for DDC. As discussed earlier, approximate dynamic programming-based methods are known to suffer from a lack of provable convergence beyond linear reward models (Jiang and Xie 2024; Tsitsiklis and Van Roy 1996a; Van Hasselt et al. 2018; Wang et al. 2021) [^4] In fact, “fitted fixed point methods can diverge even when all of the following hold: infinite data, perfect optimization with infinite computation, 1-dimensional function class (e.g., linear) with realizability (Assumption 4)(the instability is) not merely a theoretical construction: deep RL algorithms are known for their instability and training divergence…(Jiang and Xie 2024); the semi-gradient method is a popular, efficient approximation method that has theoretical assurance of convergence to projected squared Bellman error minimizers for linear reward/value functions (Sutton and Barto 2018). Feng et al. (2020) showed global concavity of value function under certain transition functions and monotonicity of value functions in terms of one-dimensional state, both of which are easily satisfied for applications in social science problems. However, those conditions are limitedly satisfied for the problems with larger dimensional state space.
@@ 2.2 Offline inverse reinforcement learning literature
The most widely used inverse reinforcement learning model, Maximum-Entropy inverse reinforcement learning (MaxEnt-IRL), assumes that the random choice happens due to agents choosing the optimal policy after penalization of the policy by its Shannon entropy (Ermon et al. 2015). In addition to the equivalence of MaxEnt-IRL to DDC (See Ermon et al. (2015) and Web Appendix $\S$3.3), the identifiability condition for DDC (Magnac and Thesmar 2002) was rediscovered by Cao et al. (2021) for MaxEnt-IRL. Zeng et al. (2023) proposes a two-step maximum likelihood-based method that can be considered as a conservative version of CCP method of Hotz and Miller (1993) [^5] When there is no uncertainty in the transition function, approximated trajectory gradient of Offline IRL method degenerates to forward simulation-based gradient in CCP estimator method of Hotz and Miller (1993).. Despite that their method is proven to be convergent, it requires the explicit estimation of the transition function, and its global convergence was only proven for linear reward functions.
Finn et al. (2016) showed that a myopic [^6] See Cao et al. (2021) for more discussion on this. version of MaxEnt-IRL can be solved by the Generative Adversarial Network (GAN) training framework (Goodfellow et al. 2020). Fu et al. (2017) extended this framework to a non-myopic version, which is proven to recover the reward function and the value function (up to policy invariance) but only under deterministic transitions. Note that GAN approaches identify rewards only up to policy invariance (Ng et al. 1999), which implies that counterfactual analysis is impossible. The GAN approach has been extended to $Q$-estimation methods that use fixed point iteration (Geng et al. 2020, 2023). Ni et al. (2021) has shown that the idea of training an adversarial network can also be used to calculate the gradient direction for minimizing the myopic version of negative log likelihood [^7] Minimizing negative log-likelihood is equivalent to minimizing KL divergence. See the Proof of Lemma 20.. Zhan et al. (2024) proposed a GAN approach that is provably equivalent to negative log likelihood minimization objective without satisfying the Bellman equation (i.e., behavioral cloning (Torabi et al. 2018)). Indeed, GAN approaches are known to work well for Imitation Learning (IL) tasks (Zare et al. 2024). However, as discussed earlier, solving for IL does not allow us to conduct a counterfactual simulation.
@@ 3.1 Setup: Maximum Entropy-Inverse Reinforcement Learning (MaxEnt-IRL)
Throughout the paper, we use $\lambda=1$, the setting which is equivalent to dynamic discrete choice (DDC) model with mean zero T1EV distribution (see Web Appendix $\S$3.3 for details); all the results of this paper easily generalize to other values of $\lambda$. Given $\pi^{*}$, we define the value function $V^{*}$ as:
@@ 3.4 Identification
[…]
@@
identifies $Q^{*}$ up to $s\in\overline{\mathcal{S}},a\in\mathcal{A}$. Furthermore, $r$ is obtained up to $\forall s\in\overline{\mathcal{S}},a\in\mathcal{A}$ by solving:
@@ 4.1 Identification via expected risk minimization
$\left\{Q\in\mathcal{Q}\mid\hat{p}_{Q}(\;\cdot\mid s)=\pi^{*}(\;\cdot\mid s)\quad\forall s\in\bar{\mathcal{S}}\quad\text{a.e.}\right\}$
@@ 4.2 Estimation via minimax-formulated empirical risk minimization
[…]
@@
[…]
@@
$\mathcal{R}_{emp}(Q;\mathcal{D}):=\frac{1}{N}\sum_{\left(s,a,s^{\prime}\right)\in\mathcal{D}}\bigl[{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}-\log\left(\hat{p}_{Q}(a\mid s)\right)}\bigr]$
$\quad\quad\quad+\frac{1}{N}\sum_{\left(s,a,s^{\prime}\right)\in\mathcal{D}}\bigl[\mathbbm{1}_{a=a_{s}}\bigl\{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bigl(\hat{\mathcal{T}}Q\left(s,a,s^{\prime}\right)-Q(s,a)\bigr)^{2}}-\beta^{2}{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\left(V_{Q}\left(s^{\prime}\right)-\bar{\zeta}(s,a)\right)^{2}}\bigr\}\bigr]$ (15)
@@ 5 GLADIUS: Algorithm for ERM-IRL (ERM-DDC)
8:  where $)V_{\bm{\theta}}(s^{\prime}):=\log\sum_{\tilde{a}\in\mathcal{A}}\exp(Q_{\bm{\theta}}(s^{\prime},\tilde{a}$))
@@
[…]
@@
[…]
When the transition function is deterministic (e.g., in Rafailov et al. (2024); Guo et al. (2025); Zhong et al. (2024)) meaning that for any state-action pair $(s,a)$, the next state $s^{\prime}$ is uniquely determined, the ascent step involving $\zeta$ is no longer required. This is because the term $\left(V_{Q}\left(s^{\prime}\right)-\zeta(s,a)\right)^{2}$ (highlighted in orange in Equation (12) and (15)) becomes redundant in the empirical ERM-IRL objective, because $\max_{\zeta\in\mathbb{R}}-\mathbb{E}_{s^{\prime}\sim P(s,a)}\left[\left(V_{Q}\left(s^{\prime}\right)-\zeta\right)^{2}\mid s,a\right]$ is always 0. Consequently, the optimization simplifies to:
@@ 6 Theory and analysis of GLADIUS
As discussed in the previous section, Equation (15) represents a mini-max optimization problem. Such problems are known to be globally solvable by a simple gradient ascent-descent algorithm when it is a concave-convex mini-max problem. However, the challenge is that Equation (12) is not a concave-convex mini-max problem. Given $Q$, it determines $\zeta$ that serves as the Bayes-optimal estimator for $\mathbb{E}_{s^{\prime}\sim P(s,a)}\left[V_{Q}\left(s^{\prime}\right)\mid s,a\right]$. This implies that $\mathbb{E}_{s^{\prime}\sim P(s,a)}\left[\left(V_{Q}\left(s^{\prime}\right)-\zeta\right)^{2}\mid s,a\right]$ is strongly concave in $\zeta$. On the other hand, given such an optimal $\zeta$, $\mathcal{L}_{BE}(Q)(s,a)$ term is not convex in $Q$ (Bas-Serrano et al. 2021). The key result in this section is proving that both $\mathcal{L}_{BE}(Q)(s,a)$ and $\mathcal{L}_{\mathrm{NLL}}(Q)(s,a)=\left[-\log\left(\hat{p}_{Q}(a\mid s)\right)\right]$, under certain parametrization of $Q$, satisfies the Polyak-Łojasiewicz (PL) condition, which allows Algorithm 1 to converge to global optima.
@@ 6.2 Global convergence of GLADIUS
Denote $\mathcal{D}_{N}$ be a finite-size dataset with $N$ number of sampled transition pairs $\left(\left(s,a,s^{\prime}\right)\right)$. Let $\Theta^{*}$ denote the (possibly non-singleton) set of expected risk (Equation (17)) minimizers. Let $\hat{\bm{\theta}}_{T}$ be the parameters returned by Algorithm 1 after $T$ SGDA iterations with step-sizes $\eta_{t}=\frac{c_{1}}{c_{2}+T}$.
@@
$\mathbb{E}_{\mathcal{D}_{N}}\left[\operatorname{dist}\left(\widehat{\bm{\theta}}_{T},\Theta^{*}\right)^{2}\right]$ $\leq O\!\left((c_{2}+T)^{-\min\left\{\tfrac{1}{2},\;\tfrac{3c\,c_{1}}{8}\right\}}\right)\;+\;\!\frac{C^{\prime}}{N}$
and
$\mathbb{E}_{\mathcal{D}_{N}}\left[\left\|Q^{\ast}-Q_{\hat{\bm{\theta}}_{T}}\right\|_{\infty}^{2}\right]\leq O\left(\left(c_{2}+T\right)^{-\min\left\{\frac{1}{2},\frac{3cc_{1}}{8}\right\}}\right)+\frac{B^{2}C^{\prime}}{N}.$ (19)
[…]
@@ 7 Offline IRL/DDC experiments
We now present results from simulation experiments, comparing the performance of our approach with that of several benchmark algorithms. In this section, we use the high-dimensional version of the canonical bus engine replacement problem (Rust (1994)) as the setting for our experiments. Later, in $\S$8, we use the OpenAI gym benchmark environment experiments with a discrete action space (Lunar Lander, Acrobot, and Cartpole) Brockman (2016) as in Garg et al. (2021) for the related but easier problem of imitation learning.
@@ 7.1 Experimental Setup
State transitions at each period.   If the manager chooses maintenance, the mileage advances by 1, 2, 3, or 4 with a $1/4$ probability each. If the manager chooses to replace the engine, then the mileage is reset to 1. That is, $\mathbb{P}(\{x_{j(h+1)}=x_{jh}+k\}\mid d_{jh}=0)=1/4$, $k\in\{1,2,3,4\}$ and $\mathbb{P}\{x_{j(h+1)}=1\mid d_{jh}=1\}=1$. When the bus reaches the maximum mileage of 20, we assume that mileage remains at 20 even if the manager continues to choose maintenance.